Helium-neon laser light of wavelength 353 nm is sent through a 0.459 mm wide single slit.

What is the width of the central maximum on a screen 1.22 m from the slit?
Answer #1

The wave length of the Helium-neon laser ? = 353nm

the width of the slit a = 0.459*10^-3 m

Then from the condition

           sin? = ?/a

                   = 353*10^-9 / 0.459*10^-3

                    = 0.000769

By uisng tangent function

           sin??tan? = y/L

then the position y = (1.22)(0.000769) = 0.000938m

Therefore the width of the central fringe

          w = 2y = 2(0.000938) = 0.001876m

              or     1.876mm